OpenStudy (anonymous):

find r and a if S_5 = 61 and a_5=81

OpenStudy (anonymous):

what does a_5 mean? and where's r in the question?

OpenStudy (anonymous):

this a geometric sequence topic :)

OpenStudy (anonymous):

ok so \(S_5\) is the sum right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[S_5=a+ar+ar^2+ar^3+ar^4\] it is the sum of the first five terms

OpenStudy (anonymous):

S_5=(A_1*rA_n)/(1-r)

OpenStudy (anonymous):

do you know the formula for adding this up?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

S_5=A_5*r^4 ?

OpenStudy (anonymous):

you can use \[\frac{a_1(1-r^n)}{1-r}\]

OpenStudy (anonymous):

so \[81=\frac{a_1-a_1r^4}{1-r}\] and also \(a_1r^4=61\) so we have \[81=\frac{a_1-61}{1-r}\] hmmm

OpenStudy (anonymous):

we use another equation

OpenStudy (anonymous):

oh what equation?

OpenStudy (anonymous):

A_5=A_1*r^4 ? so that 81=A_1*r^4 A_1=81/r^4 then substitute?

OpenStudy (anonymous):

hold on let me write it, but are you sure it is \(S_5=61\) and not \(S_5=60\)?

OpenStudy (anonymous):

yes po :)

OpenStudy (anonymous):

too bad because if it was 60 it would be easy, \(a_1=1, r=-3\)

OpenStudy (anonymous):

uhmm...can we use another eq. for S_5? :)

OpenStudy (anonymous):

lets try \[\frac{a_1-a_1r^4}{1-r}\] and plug in what we know \[\frac{a_1-81}{1-r}=60\]

OpenStudy (anonymous):

cause honestly i already solve this and i got r=1 I just dont know if thats correct just wanna check

OpenStudy (anonymous):

it cannot be right, if r = 1 it is not geometric, it is constant

OpenStudy (anonymous):

also the denominator would be undefined

OpenStudy (anonymous):

ohhkay.. :)

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

my problem is that we have one equation and two variables, but i have yet to use the fact that it is geometric what i am really thinking is that there may be infinite solutions

OpenStudy (anonymous):

lets say we do as you suggested and write \(a_1=\frac{81}{r^4}\)

OpenStudy (anonymous):

then we have \[\frac{\frac{81}{r^4}-81}{1-r}=60\] oh maybe we can solve this for \(r\)

OpenStudy (anonymous):

61 hehe

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

itll be 20r^5 + 61r^4 - 81= 0 then we use quadratic form. to find r?

OpenStudy (anonymous):

maybe my algebra is bad but that is not what i got

OpenStudy (anonymous):

wat u got?

OpenStudy (anonymous):

let me write it out

OpenStudy (anonymous):

\[\frac{\frac{81}{r^4}-81}{1-r}=61\] \[\frac{81}{r^4}-81=61-61r\] \[\frac{81}{r^4}+61r-142=0\] or \[61r^5-142r^4+81=0\]

OpenStudy (anonymous):

being a fifth degree polynomial means there is no way to solve it unless you get lucky and guess

OpenStudy (anonymous):

wat formula did u use for S_5?

OpenStudy (anonymous):

\(S_5=\frac{a_1(1-r^4)}{1-r}\)

OpenStudy (anonymous):

is that not what you are using?

OpenStudy (anonymous):

r^4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh crap yeah \(r^5\)!!

OpenStudy (anonymous):

yeah i am stuck sorry 81 looks suspiciously like \((-3)^4\) and if you use \(=1,r=-3\) unfortunately you get a sum of 61 and not 60 i have no idea how you are supposed to solve a polynomial of this degree, although there was a solution in the worlfram link i sent

OpenStudy (anonymous):

lololol you have 61!!!!

OpenStudy (anonymous):

i need coffee ok solution is done \[a=1,r=-3\] that is all and it is an easy check \[1-3+9-27+81=61\] finished!!!

OpenStudy (anonymous):

lol okay thanks i forgot we can always make our own solution :) thanks thanks :)